# Mingxuan Hu Ziqi Liu Hongbo Ma

### v1

@@ I Introduction
In recent years, automated depression detection based on artificial intelligence has gained increasing attention. Among various modalities, audio-based methods show great potential due to easy data acquisition, low cost, and non-invasiveness. Models like wav2vec 2.0 [6], HuBERT [7], and WavLM [8] have been successfully applied to learn effective speech representations, forming a strong foundation for high-performance depression detection.
@@
Regarding the issue of gender bias, researchers have proposed several mitigation strategies, most of which are pre-processing (data-level) approaches. For example, Bailey and Plumbley [12] enforced a balanced number of samples across different classes in the training set by sub-sampling the majority group. In addition, Cheong et al. [13] proposed a feature-level data augmentation technique to synthesize new virtual samples. However, these methods only adjust superficial data distribution and fail to break the spurious causal link between gender and depression within the model’s decision process. Therefore, more systematic approach is needed to eliminate such bias.
@@ III-B Causal Modeling on Depression Detection
Following common practice, we define audio-based depression detection a binary classification task, where the model predicts the depression state $d$ from the candidate set $D=\{\text{Depressed},\text{Non-depressed}\}$ using acoustic cues $C=c$ and gender information $G=g$.
@@
To achieve this goal, we compares the related factual and counterfactual scenarios, illustrated in Fig. 2LABEL:sub@fig4-sub2. In factual scenario, $G$ is in state $g$ and $C$ in state $c$, the state of $D$ can be represented as:
@@
Figure 2: (a) Csusal graph for depression detection. (b) Comparison between factual and counterfactual scenarios in depression detection. White nodes are at the state under $G=g$ and $C=c$, while gray nodes are at the state under $G=\bar{g}$ and $C=\bar{c}$.

### v2

@@ I Introduction
In recent years, automated depression detection based on artificial intelligence has gained increasing attention. Among various modalities, audio-based methods show great potential due to easy data acquisition, low cost, and non-invasiveness. Models like wav2vec 2.0 [6] have been successfully applied to learn effective speech representations, forming a strong foundation for high-performance depression detection.
@@
Regarding the issue of gender bias, researchers have proposed several mitigation strategies, most of which are pre-processing (data-level) approaches. For example, Bailey and Plumbley [10] enforced a balanced number of samples across different classes in the training set by sub-sampling the majority group. In addition, Cheong et al. [11] proposed a feature-level data augmentation technique to synthesize new virtual samples. However, these methods only adjust superficial data distribution and fail to break the spurious causal link between gender and depression within the model’s decision process. Therefore, more systematic approachs are needed to eliminate such bias.
@@ III-B Causal Modeling on Depression Detection
Following common practice, we define audio-based depression detection as a binary classification task, where the model predicts the depression state $d$ from the candidate set $D=\{\text{Depressed},\text{Non-depressed}\}$ using acoustic cues $C=c$ and gender information $G=g$.
@@
To achieve this goal, we compare the related factual and counterfactual scenarios, illustrated in Fig. 2LABEL:sub@fig4-sub2. In factual scenario, $G$ is in state $g$ and $C$ in state $c$, the state of $D$ can be represented as:
@@
Figure 2: (a) Causal graph for depression detection. (b) Comparison between factual and counterfactual scenarios in depression detection. White nodes are at the state under $G=g$ and $C=c$, while gray nodes are at the state under $G=\bar{g}$ and $C=\bar{c}$.
@@ VI Conclusion
Acknowledgment
This research was supported by National Natural Science Foundation of China (Grant No. 62502396) and XJTLU Research Development Fund (RDF-24-02-008).